### v4

# language=R,texcl=true,style=vs

Abstract
Energy statistics, also known as $\mathcal{\varepsilon}$-statistics, are functions of distances between statistical observations. This class of functions has enabled the development of non-linear statistical concepts, such as distance variance, distance covariance, and distance correlation. However, the computational burden associated with $\mathcal{\varepsilon}$-statistics is substantial, particularly when the data reside in multivariate space. To address this challenge, we have developed a method to significantly reduce memory requirements and accelerate computations, thereby facilitating the analysis of large data sets. The following cases are demonstrated: univariate and multivariate distance variance, distance covariance, partial distance correlation, energy distance, and hypothesis testing for the equality of univariate distributions.
Keywords: Median regression, multivariate median, multivariate regression
MSC: 62H11, 62H30
1 Introduction
Székely et al., (2007) pioneered the concept of energy statistics, which they termed ”energy statistics” or $\mathcal{\varepsilon}$-statistics. In their seminal work, they introduced the distance correlation (Székely and Rizzo,, 2009) as a measure of nonlinear correlation and, in essence, dependence between two random variables in arbitrary dimensions. Despite the fact that the two primary works of Székely et al., (2007), Székely and Rizzo, (2009) have amassed over 4,000 citations on Google Scholar, it could be argued that they are widely employed. However, the computational burden associated with this quantity could be perceived as a hindrance to the widespread applicability of this novel correlation, which was highly lauded by Speed, (2011) who designated it as the correlation of the 21st century.
Distance correlation has been successfully used in many fields, such as variable selection (Li et al.,, 2012), network analysis (MacCarron et al.,, 2023), and time series (Edelmann et al.,, 2019). Its uses extend beyond statistics, such as in finance (Ugwu et al.,, 2023), bioinformatics (Hou et al.,, 2022), and physics (Kasieczka and Shih,, 2020) to name a few.
Huo and Székely, (2016) proposed a fast implementation of the distance correlation, but for the univariate case. Their algorithm possesses a computational complexity class $O(nlogn)$, which is the same as required for the computation of the Spearman correlation. Their algorithm allowed for the computation of the distance correlation even with millions of observations.However, for the case of two or more dimensions, the current implementations in the package R package energy (Rizzo and Szekely,, 2024). The latter is limited by the number of observations (or sample size), and this limitation also applies to the distance variance, distance covariance, and energy distance. However, the partial distance correlation, the energy distance, and the equality of distributions tests that rely on the latter are still computationally expensive and require substantial memory when the number of observations is measured in tens of thousands or more.
The present paper makes two contributions. Firstly, it demonstrates methods for reducing the computational burden and memory requirements of certain energy statistics quantities. Secondly, it elucidates techniques for increasing the speed of calculations while reducing memory requirements. In the multivariate case, the paper employs the formulas presented in Székely and Rizzo, (2023). In the univariate case, it utilizes mathematical identities. These approaches are instrumental in expediting the calculation of energy statistics, including the energy distance, a metric of distribution equality that hinges on energy distance, distance variance, covariance, correlation, and partial distance correlation.
The next section briefly describes some quantities from the energy statistics and shows how to make computationally efficient, while Section 3 presents the relevant functions in the R package Rfast (Papadakis et al.,, 2025). Section 4 compares the Rfast implementations to implementations in other R packages and finally Section 5 concludes the paper.
2 Energy statistics
The energy distance, distance variance covariance and correlation are related to the concept of energy. Assume we have two sets of data ${\bf X},{\bf Y}\in\mathbb{R}^{p}$ [^1] The dimensions of $\bf x$ and $\bf Y$ need not be the same, but our implementation works only if they are equal..
2.1 Energy distance
The energy between ${\bf X}$ and ${\bf Y}$ is given by
$\mathcal{\varepsilon}\left({\bf X},{\bf Y}\right)=\frac{2}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\|{\bf X}_{i}-{\bf Y}_{j}\|-\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\|{\bf X}_{i}-{\bf X}_{j}\|-\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\|{\bf Y}_{i}-{\bf Y}_{j}\|,$ (1)
[…]
Similarly to the energy package, we compute the energy distance as $T=\frac{nm}{n+m}\mathcal{\varepsilon}\left({\bf X},{\bf Y}\right)$.
2.2 Distance variance, covariance and correlation
For the next three cases under study denote by $A$ and $B$ the Euclidean distance matrices of ${\bf X}$ and ${\bf Y}$, respectively, where $a_{ij}=\|{\bf X}_{i}-{\bf X}_{j}\|$ and $b_{ij}=\|{\bf Y}_{i}-{\bf Y}_{j}\|$ denote the $(i,j)$ elements of these matrices.
We next define the doubly centered matrix $\hat{A}$ whose entries are given by
$\hat{A}_{ij}=a_{ij}-\bar{a}_{i.}-\bar{a.j}+\bar{a}_{..}\ \ (i,j=1,\ldots,n),$
where
$\bar{a}_{i.}=\frac{1}{n}\sum_{j=1}^{n}a_{ij},\ \ \bar{a}_{.j}=\frac{1}{n}\sum_{i=1}^{n}a_{ij}\ \ \text{and}\ \ \bar{a}_{..}=\frac{1}{n^{2}}\sum_{i,j=1}^{n}a_{ij}.$
Similarly, the $(i,j)$ elements of the doubly centered matrix $\hat{B}$ are given by
$\hat{B}_{ij}=b_{ij}-\bar{b}_{i.}-\bar{b.j}+\bar{b}_{..}\ \ (i,j=1,\ldots,n),$
The sample distance covariance, variance, correlation and partial correlation are given by (Székely and Rizzo,, 2009, 2023)
$\mathcal{V}_{n}^{2}\left({\bf X},{\bf Y}\right)$ $=$ $\frac{1}{n^{2}}\sum_{i,j=1}\hat{A}_{ij}\hat{B}_{ij}$ (2a)
$\mathcal{V}_{n}^{2}\left({\bf X}\right)$ $=$ $\mathcal{V}_{n}^{2}\left({\bf X},{\bf X}\right)=\frac{1}{n^{2}}\sum_{i,j=1}\hat{A}_{ij}$ (2b)
$\mathcal{R}_{n}^{2}\left({\bf X},{\bf Y}\right)$ $=$ $\frac{\mathcal{V}_{n}^{2}\left({\bf X},{\bf Y}\right)}{\sqrt{\mathcal{V}_{n}^{2}\left({\bf X}\right)\mathcal{V}_{n}^{2}\left({\bf Y}\right)}}$ (2c)
$\mathcal{R}_{n}^{*2}\left({\bf X},{\bf Y}|{\bf Z}\right)$ $=$ $\frac{\mathcal{R}_{n}^{*2}\left({\bf X},{\bf Y}\right)-\mathcal{R}_{n}^{*2}\left({\bf X},{\bf Z}\right)\mathcal{R}_{n}^{*2}\left({\bf Y},{\bf Z}\right)}{\sqrt{1-\mathcal{R}_{n}^{*2}\left({\bf X},{\bf Z}\right)^{2}}\sqrt{1-\mathcal{R}_{n}^{*2}\left({\bf Y},{\bf Z}\right)^{2}}},$ (2d)
where $\mathcal{R}_{n}^{*2}\left({\bf X},{\bf Y}\right)$ denotes the unbiased distance correlation of $\bf X$ and $\bf Y$.
Let us now focus on the case of distance covariance (2a). An alternative formula is given by (Székely and Rizzo,, 2023)
$\mathcal{V}_{n}^{2}\left({\bf X},{\bf Y}\right)=\frac{\sum_{i\neq j}^{n}a_{ij}b_{ij}}{n^{2}}-\frac{2\sum_{i=1}^{n}a_{i.}b_{i.}}{n^{3}}+\frac{a_{..}b_{..}}{n^{4}}.$ (3)
The bias-corrected distance covariance entails slightly different denominators (Székely and Rizzo,, 2023)
$\mathcal{V}_{n}^{*2}\left({\bf X},{\bf Y}\right)=\frac{\sum_{i\neq j}^{n}a_{ij}b_{ij}}{n(n-3)}-\frac{2\sum_{i=1}^{n}a_{i.}b_{i.}}{n(n-2)(n-3)}+\frac{a_{..}b_{..}}{n(n-1)(n-2)(n-3)}.$ (4)
In the case of distance variance the formula changes accordingly to become
$\mathcal{V}_{n}^{2}\left({\bf X}\right)=\frac{\sum_{i\neq j}^{n}a_{ij}^{2}}{n^{2}}-\frac{2\sum_{i=1}^{n}a_{i.}^{2}}{n^{3}}+\frac{a_{..}^{2}}{n^{4}}.$ (5)
while for the unbiased distance variance the formula becomes
$\mathcal{V}_{n}^{*2}\left({\bf X}\right)=\frac{\sum_{i\neq j}^{n}a_{ij}^{2}}{n(n-3)}-\frac{2\sum_{i=1}^{n}a_{i.}^{2}}{n(n-2)(n-3)}+\frac{a_{..}^{2}}{n(n-1)(n-2)(n-3)}.$ (6)
2.3 Testing for equality of two distributions
To test for the equality of two univariate (and multivariate distributions), the test statistic is the energy distance (1) and the test statistic becomes $T=\frac{nm}{n+m}\mathcal{\varepsilon}\left({\bf x},{\bf y}\right)$. Permutations are used, where shuffling takes place between the two vectors and Eq. (1) is computed. This process is repeated $R$ times, and the proportion of times the energy distance test statistics of the permuted data exceeds the energy distance test statistic computed at the observed data serves as the approximate p-value.
2.4 Fast and light-weight computation of the distance quantities
The R packages energy (Rizzo and Szekely,, 2024) and Rfast compute the distance (co)variance using Equations (2a) and (2b), respectively. However, this approach is problematic because it necessitates the computation and storage of the distance matrix. In the case of distance covariance or distance correlation, two matrices must be computed and stored. In scenarios where the sample size of observations approaches tens of thousands, it is imperative that the computer system possesses substantial memory capacity to facilitate the requisite computations. Alternatively, one may utilize available storage space on the hard drive, though this approach only addresses the memory constraints but does not fully address the issue of computational speed.A different approach involves computing the lower triangular matrix, which requires less memory, but the resulting reduction in computational cost is only marginal.
2.4.1 Computation of distance variance, covariance, correlation and partial correlation
Rfast computes the same quantities using Equations (3) and (5), but computes the required distances ”on-the-fly”, thus never computes the full distance matrix and hence has minimum memory requirements. Let us now focus on the computation of the distance variance (5) [^2] Evidently, the same approach applies to the bias-corrected versions.. Equation (5) comprises three terms, all related to $a_{ij}$, that is, the Euclidean distance between the $i$-th and $j$-vector. The three quantities are
$\gamma_{1}=\frac{\sum_{i\neq j}^{n}a_{ij}^{2}}{n^{2}},\ \ \gamma_{2}=\frac{2\sum_{i=1}^{n}a_{i.}^{2}}{n^{3}}\ \ \text{and}\ \ \gamma_{3}=\frac{a_{..}^{2}}{n^{4}}.$ (7)
Using two nested for() loops one may compute all terms of Equation (5) at one pass. At iterations $i$ and $j$ the value of $a_{ij}$ is computed and summed. This action related to quantity $\gamma_{3}$. The value is next squared and added in two places, one place related to quantity $\gamma_{1}$ and one place related to quantity $\gamma_{2}$. The extra memory requirements, apart from the one required by the data sets themselves is the memory required by quantity $\gamma_{2}$ which must compute the sum of $a_{ij}$ with respect to $j$, thus, at the $i$-th iteration one must store the $n$-sized vector of the $a_{ij}$ values and then sum them to obtain $a_{i.}$. The same computational tricks occur for the cases of the distance covariance (3) and correlation (2c), but this time the computations involving both datasets occur simultaneously.
For the univariate distance variance, the identity in the Gini coefficient facilitates fast computation of the second term $\gamma_{2}$, of Eq. (7),
$a_{i.}=(2i-n)x_{(i)}+\sum_{i=1}^{n}x_{i}-2\sum_{k=1}^{i}x_{(k)},$
where $x_{(i)}$ denote the ordered values, in ascending order. Regarding the numerator of $\gamma_{1}$, we can do the following
$\sum_{i\neq j}^{n}a_{ij}^{2}=\sum_{i\neq j}^{n}|x_{i}-x_{j}|^{2}=n\sum_{i=1}^{n}x_{i}^{2}-\left(\sum_{i=1}^{n}x_{i}\right)^{2}.$
Since the sum of all numbers has already been computed (it is required in $\gamma_{1}$) there is no need to compute it twice, thus avoiding the computation of the sum of all pairwise squared differences. Finally, for the partial distance correlation, the implementation in R computes the necessary distance correlations and is thus computationally efficient.
2.4.2 Computation of energy distance
The energy distance (1) exploits a similar trick. The energy consists of three quantities, the sum of all pairwise distances between the matrix $\bf X$ and $\bf Y$, and the sum of all pairwise distances of $\bf X$ and $\bf Y$. The pairwise distances are computed again ”on-the-fly”, that is, each pairwise distance is computed and summed, avoiding the computation of the whole distance matrix.
2.4.3 Computations for testing the equality of two distributions
As mentioned earlier, random permutations occur repeatedly. Thus, the constant factor $\frac{nm}{n+m}$ that multiplies the energy distance $(\varepsilon({\bf x},{\bf y}))$ does not influence the p-value. With univariate distributions Eq. (1) consists of three terms
$\gamma_{1}=\sum_{i=1}^{n}\sum_{j=1}^{m}|x_{i}-y_{j}|,\ \ \gamma_{2}=\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|\ \ \text{and}\ \ \gamma_{3}=\sum_{i=1}^{m}\sum_{j=1}^{m}|y_{i}-y_{j}|.$
The use of the identity in the Gini coefficient ensures fast computation the sum of all pairwise distances,
$\gamma_{2}=\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|=2\sum_{i=1}^{n}(ix_{(i)})-(n+1)\sum_{i=1}^{n}x_{i}.$
The same formula applies to the $\gamma_{3}$ quantity that refers to the second vector. Regarding $\gamma_{1}$ this can be computed via the following identity
$\gamma_{1}=2\sum_{i=1}^{n+m}(iz_{(i)})-(n+m+1)\sum_{i=1}^{n+m}z_{i}-\gamma_{2}-\gamma_{3},$
where $z_{(1)},\ldots,z_{(n+m)}$ denotes the ordered observations of the combined samples, ${\bf z}=\left(x_{1},\ldots,x_{n},y_{1},\ldots,y_{m}\right)$. Some extra tricks take place, for instance, minimization of the repeated computations such as sums and the creation of the necessary sequences, and the use of fast functions to sample the permutations and sort the vectors.
For the case of multivariate distributions, the commands dista(x, result = "sum") and Dist(x, result = "sum") (both available in the Rfast library) compute the first term ($\gamma_{1}$) and the other two ($\gamma_{2}$ and $\gamma_{3}$), respectively. The resulting function [^3] See Appendix for an R implementation. is not faster than the available implementation in the energy library, but it is memory-saving. Note that the command Dist(x, result = "sum") can also facilitate the memory-efficient computation of the multivariate normality test (Székely and Rizzo,, 2023) as well.
Two additional tricks are a) the fact that the combined samples are sorted and b) that the total sum does not change. When permuting the data, one needs to sort the observations that fall into the first permuted sample, the rest of the observations that form the second permuted sample are already sorted. Also, instead of sorting the observations of the first permuted sample, sort their indices, since it is faster to sort integers than numeric numbers. The $\gamma_{2}$ quantity involves calculating the sum of each permuted sample. Since the total sum does not change, one only needs to compute the sum of the observations that fall within the permuted sample, the sum of the observations of the second permuted sample is easily obtained by subtraction.
3 The relevant commands in Rfast
The relevant commands in the Rfast library are edist(), dvar(), dcov(), dcor() and pdcor() and we show some examples below. First, let us generate two matrices with a few rows and a few columns [^4] The energy distance does not require the matrices to have the same number of rows..
3.1 The command edist()
The command edist() accepts two numerical matrices as its arguments and computes the energy distance (1) between the two data sets.
library(Rfast)
x <- as.matrix( iris[1:50, 1:4] )
y <- as.matrix( iris[51:100, 1:4] )
edist(x, y)
[1] 123.5538
In case the user wants to compute the energy distance matrix among three or more data sets they should assign them in a list and provide the list as a single argument in the function.
z <- as.matrix(iris[101:150, 1:4])
a <- list()
a[[ 1 ]] <- x
a[[ 2 ]] <- y
a[[ 3 ]] <- z
edist(a)
         [,1]      [,2]      [,3]
[1,]   0.0000 123.55381 195.30396
[2,] 123.5538   0.00000  38.85415
[3,] 195.3040  38.85415   0.00000
3.2 The command dvar()
The command dvar() accepts a numerical matrix and an extra logical argument (bc) whose default is FALSE and defines whether the bias-corrected distance variance (6) should be computed or not.
dvar(x)
[1] 0.2712927
dvar(x, bc = TRUE)
[1] 0.06524269
3.3 The command dcov()
The command dcov() accepts two numerical matrices and the logical argument (bc) whose default is FALSE and defines whether the bias-corrected distance covariance (4) should be returned or not.
dcov(x, y)
[1] 0.1025087
dcov(x, y, bc = TRUE)
[1] -0.002748351
3.4 The command dcor()
The command dcor() works in the same way as the command dcov but this returns a list with the (bias-corrected) distance variance of each data set, and their distance variance and correlation (2c).
dcor(x,y)
\$dcov
[1] 0.1025087
\$dvarX
[1] 0.2712927
\$dvarY
[1] 0.4135274
\$dcor
[1] 0.3060479
dcor(x, y, bc = TRUE)
\$dcov
[1] -0.002748351
\$dvarX
[1] 0.06524269
\$dvarY
[1] 0.1568211
\$dcor
[1] -0.0271709
3.5 The command pdcor()
The command pdcor() unlike dcor() returns only the partial distance correlation (2d).
pdcor(x, y, z)
[1] -0.02722611
4 Measuring the computational cost and speed-up factors
This section illustrates the time improvements comparing the functions dcor() and edit() from the packages energy, and Rfast (most recent implementation). The experiments were conducted utilising a Dell laptop equipped with Intel Core i5-1053G1 CPU at 1GHz, with 256 GB SSD, 8 GB RAM and Windows installed. Using a range of sample sizes (n=(1000, 2000, 5000, 10000, 20000, 50000)) and dimensions (p=(2, 5, 10, 20)) we generate two random matrices, and compute the running time [^5] The time was measured using the built-in command system.time(). each function requires to compute the aforementioned quantities. Table 1 presents the running time (in seconds) required by each implementation [^6] For the case of sample sizes $n=20000$ and higher, the following message appeared ”Error in .dcov(x, y, index) : ’R_Calloc’ could not allocate memory (20000 of 8 bytes)”..
The distance correlation examples were repeated, but this time compared the running time required by the function distcor() that is available in the package dcortools (Edelmann and Fiedler,, 2022) with the option to use the memory save algorithm that according to the authors of the package has a computational complexity of $O(n^{2})$ but requires only $O(n)$ memory.
4.1 Results
Table 1 and 2 presents the running times the energy and Rfast packages require for the computation of the distance correlation, energy distance, and partial distance correlation, respectively [^7] We emphasize that the implementation of the partial distance correlation is in R.. Table 3 presents the estimated scalability of the Rfast::dcor with regards to the sample size, and figure 1 visualizes Tables 1 and 2 and also contains the speed-up factor of Rfast::dcor compared to dcortools::distcor.
For the distance correlation, the speed-up factor [^8] This is the ratio of the running time required by the functions in the package energy divided by the running time required by the functions in the package Rfast. ranges from 5.5 (for the case of $n=1000$ and $p=20$) to 119 (for the case of $n=15000$ and $p=10$) with an average of 33. For the energy distance, the speed-up factor ranges from 4.6 (for the case of $n=1000$ and $p=10$) and goes up to 88 (for the case of $n=15000$ and $p=2$) with an average of 26. Finally, for the partial distance correlation, the speed-up factor ranges from 7.33 (for the case of $n=1000$ and $p=20$) up to 156 (for the case of $n=15000$ and $p=2$).
For either quantity, there are two important conclusions. The speed-up factor increases with increasing sample size and second, the implementation in the package energy cannot compute those quantities when the sample size exceeds $15000$ due to large memory requirements. This implies that, given the specifications of the computer used, installation of a larger memory that would allow the quantities to be computed with the energy package, would lead to higher speed-up factors with increasing sample sizes.
As seen in Table 3, the running time of the computation of both the the distance correlation and the energy distance via Rfast increases quadratically with respect to the sample size [^9] The same rate was estimated for the function dcortools::distcor(). Finally, Figure 1 shows that the speed-up factor of Rfast::dcor() compared to dcortools::distcor() is always greater than 1 and as the dimensionality of the data increases, so does the speed-up factor.
Finally, Table 4 contains the running time to perform the test of equality of two univariate distributions when the p-value is computed based on 999 permutations. Evidently, memory limitations forbid the application of the test with tens of thousands of observations, but this time, there is an extra limitation, that of running time. In the case of 15,000 observations, the energy::eqdist.etest requires a tremendous amount of time, which is more than a single day. In contrast, the pure R implementation requires 11 seconds for the case of 50000 observations. Hence, the speed-up factors need not be computed for this test.
Table 1: Running times (in seconds) of the energy and the Rfast implementations of the distance correlation and of the energy distance. Empty cells imply that the quantity could not be computed by the energy package implementation due to memory limitations.
Distance correlation Energy distance
Sample size Package p=2 p=5 p=10 p=20 p=2 p=5 p=10 p=20
$n=1,000$ energy 0.126 0.086 0.086 0.110 0.126 0.158 0.232 0.342
Rfast 0.016 0.008 0.009 0.020 0.003 0.017 0.026 0.017
$n=2,000$ energy 0.412 0.554 0.588 0.636 0.662 0.896 1.236 1.906
Rfast 0.040 0.050 0.051 0.062 0.021 0.028 0.041 0.081
$n=5,000$ energy 4.446 4.424 4.512 5.670 4.012 4.300 5.258 7.750
Rfast 0.275 0.308 0.342 0.404 0.137 0.172 0.424 0.731
$n=10,000$ energy 32.706 34.166 41.676 36.850 73.970 60.334 74.510 82.416
Rfast 1.166 1.334 1.472 1.764 0.538 0.798 1.825 2.912
$n=15,000$ energy 321.894 336.263 410.177 362.679 353.902 365.658 389.848 395.272
Rfast 2.803 3.212 3.451 4.227 2.006 2.369 4.130 6.710
$n=20,000$ energy - - - - - - - -
Rfast 5.079 6.008 6.189 7.540 3.289 4.330 6.984 1.954
$n=25,000$ energy - - - - - - - -
Rfast 7.911 9.324 10.371 14.227 5.072 6.599 10.209 18.004
$n=30,000$ energy - - - - - - - -
Rfast 13.681 13.770 20.622 24.061 7.152 9.425 16.454 29.571
$n=35,000$ energy - - - - - - - -
Rfast 15.481 17.854 20.228 24.587 10.422 13.830 23.301 34.683
$n=40,000$ energy - - - - - - - -
Rfast 21.725 24.420 26.223 32.256 13.611 17.417 25.143 56.997
$n=45,000$ energy - - - - - - - -
Rfast 27.491 31.672 33.408 41.153 14.088 17.775 32.716 131.637
$n=50,000$ energy - - - - - - - -
Rfast 33.959 38.181 40.899 52.201 17.105 22.413 49.346 223.218
Table 2: Running times (in seconds) of the energy and the Rfast implementations of the partial distance correlation. Empty cells imply that the quantity could not be computed by the energy package implementation due to memory limitations.
Partial Distance correlation
Sample size Package p=2 p=5 p=10 p=20
$n=1,000$ energy 0.242 0.354 0.246 0.264
Rfast 0.026 0.030 0.032 0.036
$n=2,000$ energy 1.458 0.970 1.106 1.728
Rfast 0.092 0.098 0.128 0.186
$n=5,000$ energy 11.706 12.918 12.170 13.390
Rfast 0.702 0.630 0.792 1.120
$n=10,000$ energy 264.614 235.116 238.410 277.240
Rfast 3.112 3.460 5.074 6.774
$n=15,000$ energy 1321.772 1420.980 1436.202 1271.894
Rfast 8.454 10.378 12.014 15.702
$n=20,000$ energy - - - -
Rfast 15.050 14.788 19.224 31.688
$n=25,000$ energy - - - -
Rfast 25.528 28.576 33.946 49.212
$n=30,000$ energy - - - -
Rfast 33.562 34.834 50.248 70.294
$n=35,000$ energy - - - -
Rfast 48.480 53.442 73.054 97.498
$n=40,000$ energy - - - -
Rfast 61.976 69.916 92.764 113.812
$n=45,000$ energy - - - -
Rfast 66.192 69.940 85.028 117.706
$n=50,000$ energy - - - -
Rfast 77.248 85.526 108.740 146.986
Table 3: Estimated scalability rate of the Rfast implementations of the distance correlation, partial distance correlation and of the energy distance (95% confidence interval within parentheses).
Case p=2 p=5 p=10 p=20
dcor 2.022 (1.952, 2.092) 2.131 (2.085, 2.177) 2.143 (2.078, 2.208) 2.062 (1.995, 2.130)
edist 2.210 (2.117, 2.303) 1.987 (1.852, 2.122) 2.002 (1.885, 2.119) 2.265 (2.101, 2.429)
pdcor 2.113 (2.050, 2.177) 2.117 (2.031, 2.203) 2.145 (2.060, 2.229) 2.164 (2.076, 2.253)
Refer to caption Refer to caption
(a) Distance correlation (b) Energy distance
Refer to caption Refer to caption
(c) Partial distance correlation (d) Speed-up factor of Rfast::dcor()
Figure 1: Sample size versus running time for the Rfast implementations of (a) distance correlation and (b) energy distance, (c) partial distance correlation and (d) speed-up factor of Rfast::dcor() to dcortools::distcor().
Table 4: Running times (in seconds) of the energy, Rfast implementations of the equality of univariate distributions with $999$ permutations. Empty cells imply that the quantity could not be computed by the energy package implementation due to memory limitations.
In the case of $n=15000$ the energy implementation required 516.082 seconds for only 4 permutations. Thus it would take more than 1 day to perform 999 permutations.
Sample size energy R
$n=1,000$ 5.068 0.138
$n=2,000$ 28.512 0.320
$n=5,000$ 178.040 0.782
$n=10,000$ 712.880 1.940
$n=15,000$ 516.082 3.240
$n=20,000$ 4.172
$n=25,000$ 4.734
$n=30,000$ 5.598
$n=35,000$ 7.186
$n=40,000$ 6.038
$n=45,000$ 6.706
$n=50,000$ 7.900
5 Conclusions
We showed a simple trick to reduce the memory and computational cost associated with the energy distance and the distance variance, covariance, and correlation measures. Our next goal is to transfer the code for testing the equality of univariate distributions to C++. The same is true for the partial distance correlation implementation, but in this case we also need to reduce the number of computations to make this function faster. We would also like to create more efficient functions for computing the permutation-based p-value for more hypothesis testing procedures.
Appendix
The R version of the command eqdist.etest() to test the equality of two univariate
eqdist.etest <- function(y, x, R = 999) {
    ## univariate
    #dii is the sum of all pairwise distances of x
    ni <- length(x)
    i <- 1:ni
    x <- Rfast::Sort(x)
    sx <- sum(x)
    dii <- 2 * sum(i * x) - (ni + 1) * sx
    #djj is the sum of all pairwise distances of y
    nj <- length(y)
    j <- 1:nj
    y <- Rfast::Sort(y)
    sy <- sum(y)
    s <- sx + sy
    djj <- 2 * sum(j * y) - (nj + 1) * sy
    n <- ni + nj  ## total sample size
    z <- Rfast::Sort( c(x, y) )
    ##dij <- Rfast::dista(x, y, result = "sum")
    dtot <- 2 * sum( ( 1:n ) * z ) - (n + 1) * s
    dij <- dtot - dii - djj
    stat <- dij - nj * dii / ni - ni * djj / nj
    pstat <- numeric(R)
    for ( k in 1:R ) {
      id <- Rfast::Sort.int( Rfast2::Sample.int(n, ni) )
      xp <- z[id]
      sxp <- sum(xp)
      pdii <- 2 * sum(i * xp) - (ni + 1) * sxp
      yp <- z[-id]
      syp <- s - sxp
      pdjj <- 2 * sum(j * yp) - (nj + 1) * syp
      pdij <- dtot - pdii - pdjj
      pstat[k] <- pdij - nj * pdii / ni - ni * pdjj / nj
  ( sum( pstat >= stat ) + 1 ) / (R + 1)